OpenStudy (anonymous):

Need help! Will fan and medal!! A circle is centered at (−3, 2) and has a radius of 2. Which of the following is the equation for this circle? (x + 3)2 + (y − 2)2 = 2 (x − 3)2 + (y + 2)2 = 4 (x + 3)2 + (y − 2)2 = 4 (x − 3)2 + (y + 2)2 = 2

OpenStudy (anonymous):

@prowrestler @keshaun455 @mathstudent55 @itsmichelle29 @lizzieeej @vlz6000 @campbell_st

OpenStudy (campbell_st):

well the general form of a circle is \[(x - h)^2 + (y - k)^2 = r^2\] the centre is (h, k) and r is the radius so substitute your information and get a solution.

OpenStudy (anonymous):

What numbers exactly would i be substituting @campbell_st

OpenStudy (campbell_st):

well you are told the centre is (-3, 2) and radius is 2 so h = -3 k = 2 and r = 2

OpenStudy (anonymous):

soo (x-3)2+ (y-2)2=4? @campbell_st

OpenStudy (campbell_st):

no (x - (-3))^2 + (y -2)^2 = 2^2 simplify the 1st term

OpenStudy (anonymous):

What 1st term?? I'm so confused @campbell_st

OpenStudy (anonymous):

Would the answer be B then?

OpenStudy (anonymous):

@k_lynn @campbell_st

OpenStudy (campbell_st):

the answer isn't B.. just substitute correctly h = - 3 so (x - h) = ( x -(-3) which becomes (x + 3)